OpenStudy (anonymous):

find the measures of 2 supplementary angles if one angle is 12 more than 3 times the other find the measures of 2 supplementary angles if one angle is 12 more than 3 times the other @Mathematics

OpenStudy (anonymous):

Let one angle be x. The other angle is 12 + 3x. Since they are supplementary, 12 + 3x + x = 180 12 + 4x = 180 4x = 180 -12 x = 168 / 4 x = 42 Hence the two angles are 42 degrees and (180 - 42) = 138 degrees.

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

no problem :)